# The formation of a soliton gas condensate for the focusing Nonlinear Schrödinger equation

### v1

@@ Abstract
In this work, we carry out a rigorous analysis of a multi-soliton solution of the focusing nonlinear Schrödinger equation as the number, $N$, of solitons grows to infinity. We discover configurations of $N$-soliton solutions which exhibit the formation (as $N\to\infty$) of a soliton gas condensate. Specifically, we show that when the eigenvalues of the Zakharov - Shabat operator for the NLS equation accumulate on two bounded horizontal segments in the complex plane with norming constants bounded away from 00, then, asymptotically, the solution is described by a rapidly oscillatory elliptic-wave with constant velocity, on compact subsets of $(x,t)$. This is to be distinguished from previous analyses of soliton gasses where the norming constants were tending to zero with $N$, and the asymptotic description only included elliptic waves in the long-time asymptotics.
@@ 1 Introduction
The soliton, and $N$-soliton solutions, were part of the original discovery of integrable nonlinear partial differential equations, and there emerged very quickly a dual interpretation of the “solitonic component” of a solution as a collection of particles, at least in the $|t|\to\infty$ asymptotic regime. Of course, for intermediate times, it is not easy to identify any individual solitons in a solution of the PDE, but nonetheless the particle interpretation persisted, with Zakharov [40] proposing a kinetic theory of solitons in 1971. The kinetic theory of solitons is a qualitative theory describing the evolution of a density function for solitons, supposing them to be individual particles which asymptotically interact in a pair-wise fashion. In the last decade, there has emerged new experimental, numerical, and analytical justification for this qualitative theory (see [36] for a review of these developments).
In this paper we develop the analytical description of $N$-soliton solutions in a new region of space-time which has been numerically observed (and conjectured) to correspond to a special type of soliton gas called a soliton condensate. The paper is organized as follows. In Section 2 we present a very brief introduction to the kinetic theory of solitons, explain the characterization of $N$-soliton solutions in the soliton gas condensate scaling via Riemann-Hilbert problems, formulate a Riemann-Hilbert problem for the limiting soliton gas condensate solution, and describe our results regarding the behavior of $N$-soliton solutions of the NLS equation in the condensate scaling. In Section 3-6 we carry out the asymptotic analysis of the Riemann-Hilbert problem for the soliton gas condensate. In Section 7 we present the proof of our main results, using the previously mentioned asymptotic analysis. Some technical results are deferred to the appendices.
@@ 2.1 Direct and Inverse Scattering for the NLS equation
The transmission coefficient also has an analytic extension to $\mathbb{C}_{+}$, and each pole of $T(z)$ in $\mathbb{C}_{+}$ corresponds to an $L^{2}$ eigenfunction for the differential equation (1.2). For reasonable classes of functions $\psi(x)$, there are at most a finite number of poles of $T(z)$ in $\mathbb{C}_{+}$. At each pole $\lambda_{j}$, the first column of $\Phi^{-}(x,\lambda_{j})$ and the second column of $\Phi^{+}(x,\lambda_{j})$ are linearly dependent, and the proportionality constant $c_{j}$ relating them is referred to as a normalization constant, and is defined by
@@
Given spectral data $\mathcal{D}(t)$ find a $2\times 2$ matrix-valued function $M(z)$ such that
@@
    $\text{res}_{\lambda_{j}}M(z)=\lim_{z\to\lambda_{j}}\left[M(z)\begin{pmatrix}0&0\\ c_{j}e^{2i\theta(\lambda_{j};x,t)}&0\end{pmatrix}\right]\,,$ (2.7)
@@
    $\text{res}_{{\overline{\lambda_{j}}}}M(z)=\lim_{z\to{\overline{\lambda}}_{j}}\left[M(z)\begin{pmatrix}0&-{\overline{c}}_{j}e^{-2i\theta({\overline{\lambda}}_{j};x,t)}\\ 0&0\end{pmatrix}\right].$ (2.8)
@@
Existence and uniqueness for this Riemann-Hilbert problem follows from what are by now standard arguments: uniqueness follows from Liouville’s theorem, and existence follows from a careful application of a vanishing lemma as described in [42]. The solution $\psi(x,t)$ to the NLS equation corresponding to the initial data $\psi(x)$ and determined by the scattering data $\mathcal{D}(t)$ is obtained from $M(z;x,t)$ through the $z\to\infty$ normalization:
@@ 2.2 $N$-soliton solutions and soliton gasses
One experimental prediction in [40] is that a tracer soliton with arbitrary eigenvalue parameter, launched into this gas, should interact with the gas in such away that its velocity deforms as it propagates, and the explicit computation of this time-dependent velocity, as a function of density of the gas as well as the tracer’s initial eigenvalue parameter, yielded an experimental prediction which could in principle be tested. About 30 years later, El [16] extended the kinetic theory to a dense gas, in which a coupled system of equations was derived, with the evolving density coupled to the evolving velocity of a tracer. Both [40] and [16] focused on the KdV equation, but this was extended by El and Kamchatnov in [13] to other integrable nonlinear partial differential equations, including the NLS equation. Significant progress regarding the understanding of the kinetic equations arising in different settings, as well as more detailed analysis of the kinetic equations themselves, has continued, for example in [14, 27, 37, 38].
@@ 2.3 $N$-soliton solutions: condensate scaling
We will refer to such an $N$-soliton solution, or a sequence of such $N$-soliton solutions (indexed by $N$), as $N$-soliton condensate(s), and we will refer to the associated scattering data as $N$-soliton condensate scattering data. Specifically, we consider the following assumptions.
@@
    There is a density function $\rho(z)$ which is analytic in a neighborhood of the line segment $\eta_{1}$ (and the neighborhood’s Schwarz reflection), satisfying
    $\rho(z)\mbox{is analytic in a neighborhood of each line segment }\eta_{1}\mbox{ and }\eta_{2},$ (2.15)
    $\rho(z)\in{\mathbb{R}}\mbox{ for }z=t+ib,\ t\in(-a,a),$ (2.16)
@@
    There is a “normalization constant sampling function” $h(z)$ which is analytic in a neighborhood of the line segment $\eta_{1}$ (and the neighborhood’s Schwarz reflection), satisfying $h({\overline{z}})={\overline{h(z)}}$.
@@
where $\lambda_{j}$ is defined in (2.19) and the normalzation constants $c_{j}$ are defined in (2.20). For each positive integer $N$, this scattering data corresponds to an $N$-soliton condensate solution to the NLS equation, which we will denote by
@@ 2.4 Riemann-Hilbert problem for $\psi_{N}(x,t)$
Given the scattering data $\mathcal{D}_{N}$, the Riemann-Hilbert problem 2.1 simplifies significantly, since $r(z)\equiv 0$ for all real $z$. Indeed, since the jump relation across the real axis (2.6) becomes the trivial relation $M_{+}(z)=M_{-}(z)$ for all real $z$, Morera’s Theorem tells us that $M(z)$ is then analytic across the real axis, and hence $M(z)$ is meromorphic in $\mathbb{C}$. We will make one standard transformation, from $M(z)$ to a new matrix $A(z)$, following (for example) [17]. We use the contours $\Gamma_{1}$ and $\Gamma_{2}$ which are provided to us under the $N$-soliton condensate scattering assumption, $\Gamma_{1}$ encircling the interval $\eta_{1}$ and $\Gamma_{2}$ its Schwarz reflection. Each of these contours is taken to be oriented clockwise, as shown in Fig. 2. We then define $A(z)$ as follows:
@@
$A(z)=\begin{cases}M(z)\begin{pmatrix}1&0\\ -\sum_{j=1}^{N}\frac{c_{j}}{z-\lambda_{j}}e^{2i\theta(\lambda_{j})}&1\end{pmatrix},&z\textit{ inside }\Gamma_{1}\\ \\[2.0pt] M(z)\begin{pmatrix}1&\sum_{j=1}^{N}\frac{{\overline{c}}_{j}}{z-{\overline{\lambda}}_{j}}e^{-2i\theta({\overline{\lambda}}_{j})}\\ 0&1\end{pmatrix},&z\textit{ inside }\Gamma_{2}\\ \\[2.0pt] M(z),&\textit{otherwise}.\end{cases}$ (2.23)
@@
Now $M(z)$ has simple poles at the eigenvalues $\lambda_{j}$ and $\overline{\lambda_{j}}$, but the definition above is chosen so that $A(z)$ is analytic inside the contours $\Gamma_{1}$ and $\Gamma_{2}$. Indeed, the reader may check from the residue conditions (2.7) and (2.8) that (for example) the pole of $M_{11}(z)$ at $\lambda_{k}$ is cancelled since
$A_{11}(z)=M_{11}(z)-\frac{c_{k}M_{12}(z)e^{2i\theta(\lambda_{k})}}{z-\lambda_{k}}+\left<\mbox{analytic at $\lambda_{k}$}\right>,$ (2.24)
@@
and the first two terms have the same exact residue at $\lambda_{k}$. Similar arguments show that the matrix $A(z)$ is analytic inside $\Gamma_{1}$ and inside $\Gamma_{2}$, and it is also obviously analytic for $z$ outside the two contours. Finally, $A$ clearly has boundary values as $z$ approaches the contours $\Gamma_{1}$ and $\Gamma_{2}$ (with different boundary values as $z$ approaches each contour from inside or outside).
@@
Notation: boundary values for $z$ on a contour: Following standard convention, for $z$ in the contour $\Gamma_{1}$ (or $\Gamma_{2}$), we will let $A_{+}(z)$ denote the boundary value of $A$ obtained by letting $z^{\prime}$ approach $z$ from outside $\Gamma_{1}$ (or outside $\Gamma_{2}$):
$A_{+}(z)=\lim_{\begin{subarray}{c}z^{\prime}\to z\\ z^{\prime}\in\mbox{ \small exterior of }\Gamma_{j}\end{subarray}}A(z^{\prime}),\ \ z\in\Gamma_{j},$ (2.25)
@@
and we will let $A_{-}(z)$ denote the boundary value from inside $\Gamma_{1}$ (or $\Gamma_{2}$). (Note that this is consistent with the usual convention that the $+$ side of a contour is on the left of the contour as one traverses it according to its orientation).
@@
The new matrix-valued function $A(z)$ satisfies a Riemann-Hilbert problem, which we will refer to as the $N$-soliton condensate Riemann-Hilbert problem.
@@
Find a $2\times 2$ matrix function $A=A(z)=A(z;x,t,N)$ such that
  • $A$ is analytic on $\mathbb{C}\setminus\{\Gamma_{1}\cup\Gamma_{2}\}$
  • $A_{+}(z)=A_{-}(z)J_{A}(z)$ where the jump matrix $J_{A}(z)$ is given by
    $J_{A}(z)=\begin{cases}\begin{pmatrix}1&0\\ \sum_{j=1}^{N}\frac{c_{j}}{z-\lambda_{j}}e^{2i\theta(\lambda_{j})}&1\end{pmatrix},\quad&z\in\Gamma_{1}\\[4.0pt] \\ \begin{pmatrix}1&-\sum_{j=1}^{N}\frac{{\overline{c}}_{j}}{z-{\overline{\lambda}}_{j}}e^{-2i\theta({\overline{\lambda}}_{j})}\\ 0&1\end{pmatrix},\quad&z\in\Gamma_{2}\end{cases}$ (2.27)
  • $A(z)=I+\frac{1}{2iz}\left(\begin{array}[]{cc}-\int_{x}^{\infty}|\psi_{N}(s,t)|^{2}{\rm d}s&\psi_{N}(x,t)\\ \overline{\psi_{N}(x,t)}&\int_{x}^{\infty}|\psi_{N}(s,t)|^{2}{\rm d}s\end{array}\right)+\mathcal{O}\left(\frac{1}{z^{2}}\right),\ \mbox{ as }z\to\infty$.
Note that the above $z\to\infty$ asymptotic condition describes exactly how to extract the $N$-soliton condensate solution $\psi_{N}(x,t)$ from the solution $A$ of the Riemann-Hilbert problem. Usually the asymptotic normalization condition is described in the weak form $A=I+\mathcal{O}\left(\frac{1}{z}\right)$, but in this case, since the contours $\Gamma_{1}$ and $\Gamma_{2}$ are bounded, $A$ possesses a complete asymptotic expansion in powers of $\frac{1}{z}$.

### v2

@@ Abstract
In this work, we carry out a rigorous analysis of a multi-soliton solution of the focusing nonlinear Schrödinger equation as the number, $N$, of solitons grows to infinity. We discover configurations of $N$-soliton solutions which exhibit the formation (as $N\to\infty$) of a soliton gas condensate. Specifically, we show that when the eigenvalues of the Zakharov - Shabat operator for the NLS equation accumulate on two bounded horizontal segments in the complex plane with norming constants bounded away from 00, then, asymptotically, the solution is described by a rapidly oscillatory elliptic-wave with constant velocity, on compact subsets of $(x,t)$. We then consider more complex solutions with an extra soliton component, and provide rigorous justification of the predictions of the kinetic theory of solitons in this deterministic setting. This is to be distinguished from previous analyses of soliton gasses where the norming constants were tending to zero with $N$, and the asymptotic description only included elliptic waves in the long-time asymptotics.
@@ 1 Introduction
The soliton, and $N$-soliton solutions, were part of the original discovery of integrable nonlinear partial differential equations, and there emerged very quickly a dual interpretation of the “solitonic component” of a solution as a collection of particles, at least in the $|t|\to\infty$ asymptotic regime. Of course, for intermediate times, it is not easy to identify any individual solitons in a solution of the PDE, but nonetheless the particle interpretation persisted, with Zakharov [41] proposing a kinetic theory of solitons in 1971. The kinetic theory of solitons is a qualitative theory describing the evolution of a density function for solitons, supposing them to be individual particles which asymptotically interact in a pair-wise fashion. The density function is coupled to a "velocity profile function" corresponding to the effective velocity of a single particle as it interacts with the gas. In the last decade, there has emerged new experimental, numerical, and analytical justification for this qualitative theory (see [37] for a review of these developments).
In this paper we develop the analytical description of $N$-soliton solutions in a new region of space-time which has been numerically observed (and conjectured) to correspond to a special type of soliton gas called a soliton condensate. The paper is organized as follows. In Section 2 we present a very brief introduction to the kinetic theory of solitons, explain the characterization of $N$-soliton solutions in the soliton gas condensate scaling via Riemann-Hilbert problems, formulate a Riemann-Hilbert problem for the limiting soliton gas condensate solution, and describe our results regarding the behavior of $N$-soliton solutions of the NLS equation in the condensate scaling. In Section 3-6 we carry out the asymptotic analysis of the Riemann-Hilbert problem for the soliton gas condensate. In Section 7 we present the proof of our main results, using the previously mentioned asymptotic analysis. In the brief Section 8 we consider the interaction of a tracer soliton with the gas, and show that the kinetic theory of solitons holds for this soliton condensate. Some technical results are deferred to the appendices.
@@ 2.1 Direct and Inverse Scattering for the NLS equation
The transmission coefficient also has a meromorphic extension to $\mathbb{C}_{+}$, and each pole of $T(z)$ in $\mathbb{C}_{+}$ corresponds to an $L^{2}$ eigenfunction for the differential equation (1.2). For functions $\psi(x)$ in an open dense subset of $L^{1}({\mathbb{R}})$, there are at most a finite number of poles of $T(z)$ in $\mathbb{C}_{+}$. At each pole $\lambda_{j}$, the first column of $\Phi^{-}(x,\lambda_{j})$ and the second column of $\Phi^{+}(x,\lambda_{j})$ are linearly dependent, and the proportionality constant $c_{j}$ relating them is referred to as a normalization constant, and is defined by
@@
Given spectral data $\mathcal{D}(t)$, find a $2\times 2$ matrix-valued function $M(z)$ such that
@@
    $\operatorname*{\textrm{res}}_{\lambda_{j}}M(z)=\lim_{z\to\lambda_{j}}\left[M(z)\begin{pmatrix}0&0\\ c_{j}e^{2i\theta(\lambda_{j};x,t)}&0\end{pmatrix}\right]\,,$ (2.7)
@@
    $\operatorname*{\textrm{res}}_{{\overline{\lambda_{j}}}}M(z)=\lim_{z\to{\overline{\lambda}}_{j}}\left[M(z)\begin{pmatrix}0&-{\overline{c}}_{j}e^{-2i\theta({\overline{\lambda}}_{j};x,t)}\\ 0&0\end{pmatrix}\right].$ (2.8)
@@
Existence and uniqueness for this Riemann-Hilbert problem follows from what are by now standard arguments: uniqueness follows from Liouville’s theorem, and existence follows from a careful application of a vanishing lemma as described in [43]. The solution $\psi(x,t)$ to the NLS equation corresponding to the initial data $\psi(x)$, assuming that $\psi,\psi_{x}\in L^{1}({\mathbb{R}})$, and determined by the scattering data $\mathcal{D}(t)$ is obtained from $M(z;x,t)$ through the $z\to\infty$ normalization:
@@ 2.2 $N$-soliton solutions and soliton gasses
One experimental prediction in [41] is that a “tracer soliton”, a soliton whose amplitude is much bigger than all the other solitons, launched into this gas, should interact with the gas in such away that its velocity deforms as it propagates, and the explicit computation of this time-dependent velocity, as a function of density of the gas as well as the tracer’s initial eigenvalue parameter, yielded an experimental prediction which could in principle be tested. About 30 years later, El [18] extended the kinetic theory to a dense gas, in which a coupled system of equations was derived, with the evolving density coupled to the evolving velocity of a tracer. Both [41] and [18] focused on the KdV equation, but this was extended by El and Kamchatnov in [14] to other integrable nonlinear partial differential equations, including the NLS equation. Significant progress regarding the understanding of the kinetic equations arising in different settings, as well as more detailed analysis of the kinetic equations themselves, has continued, for example in [15, 28, 38, 39].
@@ 2.3 $N$-soliton solutions: condensate scaling
We will refer to the sequence of such $N$-soliton solutions as an $N$-soliton condensate, and we will refer to the associated sequence scattering data as $N$-soliton condensate scattering data. Specifically, we consider the following assumptions.
@@
    There is a density function $\rho(z)$ which is analytic in a neighborhood ${\mathcal{U}}$ of the domain enclosed by the line segment $\eta_{1}$ and the arc of the circle of radius $|A|$ centered at 00 connecting $A$ with $\overline{A}$ ($\rho$ is also assumed to be analytic in the Schwarz reflection of ${\mathcal{U}}$). The density $\rho$ satisfies
    $\rho(z)>0\mbox{ for }z=t+ib,\ t\in(-a,a),$ (2.15)
@@
    There is a “normalization constant sampling function” $h(z)$ which is analytic in ${\mathcal{U}}$ and $\overline{{\mathcal{U}}}$ satisfying $h({\overline{z}})={\overline{h(z)}}$, and $h(z)>0$ for $z\in\eta_{1}\cup\eta_{2}$.
@@
where $\lambda_{j}$ is defined in (2.18) and the normalzation constants $c_{j}$ are defined in (2.19). For each positive integer $N$, these scattering data correspond to an $N$-soliton solution to the NLS equation, which we will denote by
@@ 2.4 Riemann-Hilbert problem for $\psi_{N}(x,t)$
Given the scattering data $\mathcal{D}_{N}$, the Riemann-Hilbert problem 2.1 simplifies significantly, since $r(z)\equiv 0$ for all real $z$. Indeed, since the jump relation across the real axis (2.6) becomes the trivial relation $M_{+}(z)=M_{-}(z)$ for all real $z$, Morera’s Theorem tells us that $M(z)$ is then analytic across the real axis, and hence $M(z)$ is meromorphic in $\mathbb{C}$. We will make one standard transformation, from $M(z)$ to a new matrix $A(z)$, following (for example) [19]. We use the contours $\Gamma_{1}$ and $\Gamma_{2}$ which are provided to us under the $N$-soliton condensate scattering assumption, $\Gamma_{1}$ encircling the interval $\eta_{1}$ and $\Gamma_{2}$ its Schwarz reflection. Each of these contours is taken to be oriented clockwise, as shown in Fig. 2. We then define $A_{N}(z)$ as follows:
@@
$A_{N}(z)=\begin{cases}M(z)\begin{pmatrix}1&0\\ -\sum_{j=1}^{N}\frac{c_{j}}{z-\lambda_{j}}e^{2i\theta(\lambda_{j})}&1\end{pmatrix},&z\textit{ inside }\Gamma_{1}\\ \\[2.0pt] M(z)\begin{pmatrix}1&\sum_{j=1}^{N}\frac{{\overline{c}}_{j}}{z-{\overline{\lambda}}_{j}}e^{-2i\theta({\overline{\lambda}}_{j})}\\ 0&1\end{pmatrix},&z\textit{ inside }\Gamma_{2}\\ \\[2.0pt] M(z),&\textit{otherwise}.\end{cases}$ (2.22)
@@
Now $M(z)$ has simple poles at the eigenvalues $\lambda_{j}$ and $\overline{\lambda_{j}}$, but the definition above is chosen so that $A_{N}(z)$ is analytic inside the contours $\Gamma_{1}$ and $\Gamma_{2}$. Indeed, the reader may check from the residue conditions (8.1) and (8.2) that (for example) the pole of $M_{11}(z)$ at $\lambda_{k}$ is canceled since
$A_{N;11}(z)=M_{11}(z)-\frac{c_{k}M_{12}(z)e^{2i\theta(\lambda_{k})}}{z-\lambda_{k}}+\left<\mbox{analytic at $\lambda_{k}$}\right>,$ (2.23)
@@
and the first two terms have the same exact residue at $\lambda_{k}$. Similar arguments show that the matrix $A_{N}(z)$ is analytic inside $\Gamma_{1}$ and inside $\Gamma_{2}$, and it is also obviously analytic for $z$ outside the two contours. Finally, $A_{N}(z)$ clearly has boundary values as $z$ approaches the contours $\Gamma_{1}$ and $\Gamma_{2}$ (with different boundary values as $z$ approaches each contour from inside or outside).
@@
Notation: boundary values for $z$ on a contour: Following standard convention, for $z$ in the contour $\Gamma_{1}$ (or $\Gamma_{2}$), we will let $A_{N,+}(z)$ denote the boundary value of $A_{N}(z)$ obtained by letting $z^{\prime}$ approach $z$ from outside $\Gamma_{1}$ (or outside $\Gamma_{2}$):
$A_{N,+}(z)=\lim_{\begin{subarray}{c}z^{\prime}\to z\\ z^{\prime}\in\mbox{ \small exterior of }\Gamma_{j}\end{subarray}}A_{N}(z^{\prime}),\ \ z\in\Gamma_{j},$ (2.24)
@@
and we will let $A_{N,-}(z)$ denote the boundary value from inside $\Gamma_{1}$ (or $\Gamma_{2}$). (Note that this is consistent with the usual convention that the $+$ side of a contour is on the left of the contour as one traverses it according to its orientation).
@@
The new matrix-valued function $A_{N}(z)$ satisfies a Riemann-Hilbert problem, which we will refer to as the $N$-soliton condensate Riemann-Hilbert problem.
@@
Find a $2\times 2$ matrix function $A_{N}=A_{N}(z)=A_{N}(z;x,t)$ such that
  • $A_{N}$ is analytic on $\mathbb{C}\setminus\{\Gamma_{1}\cup\Gamma_{2}\}$
  • $A_{N,+}(z)=A_{N,-}(z)J_{A}(z)$ where the jump matrix $J_{A_{N}}(z)$ is given by
    $J_{A_{N}}(z)=\begin{cases}\begin{pmatrix}1&0\\ \sum_{j=1}^{N}\frac{c_{j}}{z-\lambda_{j}}e^{2i\theta(\lambda_{j})}&1\end{pmatrix},\quad&z\in\Gamma_{1}\\[4.0pt] \\ \begin{pmatrix}1&-\sum_{j=1}^{N}\frac{{\overline{c}}_{j}}{z-{\overline{\lambda}}_{j}}e^{-2i\theta({\overline{\lambda}}_{j})}\\ 0&1\end{pmatrix},\quad&z\in\Gamma_{2}\end{cases}$ (2.26)
  • $A_{N}(z)=I+\frac{1}{2iz}\left(\begin{array}[]{cc}-\int_{x}^{\infty}|\psi_{N}(s,t)|^{2}{\rm d}s&\psi_{N}(x,t)\\ \overline{\psi_{N}(x,t)}&\int_{x}^{\infty}|\psi_{N}(s,t)|^{2}{\rm d}s\end{array}\right)+\mathcal{O}\left(\frac{1}{z^{2}}\right),\ \mbox{ as }z\to\infty$.
Note that the above $z\to\infty$ asymptotic condition describes exactly how to extract the $N$-soliton condensate solution $\psi_{N}(x,t)$ from the solution $A_{N}$ of the Riemann-Hilbert problem. Usually the asymptotic normalization condition is described in the weak form $A_{N}=I+\mathcal{O}\left(\frac{1}{z}\right)$, but in this case, since the contours $\Gamma_{1}$ and $\Gamma_{2}$ are bounded, $A$ possesses a complete asymptotic expansion in powers of $\frac{1}{z}$.